# NNLO Global Analysis of Polarized Parton Distribution Functions

### v1

@@ Abstract
where the terms on the right-hand-side successively provide the leading order (LO), next-to-leading order (NLO), and NNLO approximations to the kernels. The NNLO terms have been computed in Refs. [7, 8, 9, 10, 11], and implemented in a modified version of the PEGASUS evolution library [12].
@@
For our NLO and NNLO global fits we adopt the same flexible functional form as in [32] for the helicity parton densities at the initial scale $Q_{0}=1$ GeV, for instance,
@@
with free parameters $N_{i}$, $\alpha_{i}$, $\beta_{i}$, $\gamma_{i}$, $\eta_{i}$, and $\kappa_{i}$. We enforce positivity $|\Delta f|/f\leq 1$ of the parton densities, using the unpolarized distributions $f(x,Q^{2})$ of [34], from which we also adopt the running of the strong coupling at each corresponding order, and which we also use for the spin-averaged cross sections in the denominators of the measured spin asymmetries. We work in the zero-mass variable-flavor-number scheme, including the corresponding matching coefficients at threshold from [35] for the charm and bottom distributions.
For processes involving observed hadrons in the final state, we need a set of fragmentation functions (FFs). As no full global analysis of FFs is available at NNLO accuracy [^3] NNLO analyses of FFs using only $e^{+}e^{-}$ and SIDIS data do exist [36, 37]. Unfortunately, these data sets do not allow for a proper extraction of the gluon FF, which is needed for $pp\to hX$., we have to rely on the most comprehensive NLO fits from [38, 39]. While some effects of the FFs cancel in the spin asymmetries, this choice introduces an extra source of theoretical uncertainty in our NNLO result. Additional challenges arise in the description of the available SIDIS data at $x\lesssim 0.1$, where the associated values of $Q^{2}$ are below the regime of validity of the unpolarized baseline PDFs and where SIDIS is not well described by current FFs [38]. Furthermore, in this regime the full NNLO corrections to SIDIS appear to be especially sizable [17] and are not represented well by the approximate NNLO results. Indeed, we have found that fits that include the SIDIS spin asymmetry data at $x\lesssim 0.1$ do not work particularly well and show a significant decrease in quality when going from NLO to NNLO, consistent with observations made in Ref. [6]. Therefore, in order to be conservative and to ensure a robust analysis, we apply a cut $x>0.12$ to the SIDIS data included in the fit, which results in comparable $\chi^{2}$-values for SIDIS at NLO and NNLO accuracy in our analysis. Commensurate with this cut, we also apply a cut of $p_{T}>1.5$ GeV for the RHIC $pp\rightarrow hX$ data.
@@
A remarkable perturbative stability is observed, with minimal changes of the PDFs from NLO to NNLO. This is also reflected in the total $\chi^{2}$–values of the fit after application of the cuts, which are reported in Table 1. The inclusion of NNLO corrections leads to improvements, albeit small, across all sets of data with the exception of $pp\to hX$, where $\chi^{2}$ remains essentially unchanged.
@@
As expected, the total up and down quark PDFs are well determined. The gluon helicity PDF is seen to be clearly positive, with relatively small uncertainty, over a broad region in $x$ accessed by RHIC data. Furthermore, there is a pronounced pattern of positive $\Delta\bar{u}$ and negative $\Delta\bar{d}$ at moderately large $x$, while the strangeness helicity PDF is essentially undetermined. This is in contrast to our previous analyses [32, 2], and results from the fact that we no longer impose constraints related to SU(2) and SU(3) symmetry [^4] Nevertheless, the central NLO and NNLO fits both deviate only by a few per cent from these constraints that are usually given in terms of $F$ and $D$ values.. Sizable uncertainties remain for the sea quark and gluon PDFs at $x\lesssim 10^{-3}$, consistent with the lack of experimental data constraining that region.
Figures 35 show detailed comparisons of our fit results at NLO and NNLO accuracy to a few selected sets of $pp$ and SIDIS data that were used in our analysis. For comparison, we also show in Figs. 15 the results for an (unpublished) NLO fit (labeled as BDSSV22) that was obtained based on a fixed flavor number scheme, without the cuts on the SIDIS data and by making use of the RHIC dijet data (for a collection of RHIC data, see, e.g. [40]). As can be seen, the differences to our new NLO fit are rather moderate and mostly within the estimated uncertainty bands.
@@
Our study sets the stage for future NNLO global analyses of nucleon helicity structure in the EIC era. There are several avenues for future improvements, among them derivations of better approximations for the NNLO terms in $pp$ scattering, the implementation of the full NNLO results for SIDIS and detailed explorations of the residual scale dependence and power corrections.

### v3

@@ Abstract
where the terms on the right-hand-side successively provide the leading order (LO), next-to-leading order (NLO), and NNLO approximations to the kernels. The NNLO terms have been computed in Refs. [7, 8, 9, 10, 11], and implemented in a modified version of the PEGASUS evolution library [12]. The evolution of polarized PDFs up to NNLO was benchmarked against the evolution libraries EKO [13] and APFEL [14], finding excellent agreement.
@@
For our NLO and NNLO global fits we adopt the same flexible functional form as in [34] for the helicity parton densities of light-quarks and gluon at the initial scale $Q_{0}=1$ GeV, for instance,
@@
with free parameters $N_{i}$, $\alpha_{i}$, $\beta_{i}$, $\gamma_{i}$, $\eta_{i}$, and $\kappa_{i}$. We enforce positivity $|\Delta f|/f\leq 1$ of the parton densities, using the unpolarized distributions $f(x,Q^{2})$ of [36], from which we also adopt the running of the strong coupling at each corresponding order, and which we also use for the spin-averaged cross sections in the denominators of the measured spin asymmetries. The distributions corresponding to heavy-quarks are generated perturbatively above their mass thresholds in the zero-mass variable-flavor-number scheme, including the corresponding matching coefficients at threshold from [37]. We do not consider heavy-quark masses effects, which are expected to be small compared to the current experimental uncertainties, although they will play a more prominent role in the future EIC [38].
For processes involving observed hadrons in the final state, we need a set of fragmentation functions (FFs). As no full global analysis of FFs is available at NNLO accuracy [^3] NNLO analyses of FFs using only $e^{+}e^{-}$ and SIDIS data do exist [39, 40]. Unfortunately, these data sets do not allow for a proper extraction of the gluon FF, which is essential for the theoretical description of $pp\to hX$ results., we have to rely on the most comprehensive NLO fits from [41, 42]. While some effects of the FFs cancel in the spin asymmetries, this choice introduces an extra source of theoretical uncertainty in our NNLO result. Additional challenges arise in the description of the available SIDIS data at $x\lesssim 0.1$, where the associated values of $Q^{2}$ are below the regime of validity of the unpolarized baseline PDFs and where SIDIS is not well described by current FFs [41]. Furthermore, in this regime the full NNLO corrections to SIDIS appear to be especially sizable [19] and are not represented well by the approximate NNLO results. Indeed, we have found that fits that include the SIDIS spin asymmetry data at $x\lesssim 0.1$ do not work particularly well and show a significant decrease in quality when going from NLO to NNLO, consistent with observations made in Ref. [6]. Therefore, in order to be conservative and to ensure a robust analysis, we apply a cut $x>0.12$ to the SIDIS data included in the fit, which results in comparable $\chi^{2}$-values for SIDIS at NLO and NNLO accuracy in our analysis. Commensurate with this cut, we also apply a cut of $p_{T}>1.5$ GeV for the RHIC $pp\rightarrow hX$ data.
@@
A remarkable perturbative stability is observed, with minimal changes of the PDFs from NLO to NNLO both in terms of their central values and uncertainties. This is also reflected in the total $\chi^{2}$–values of the fit after application of the cuts, which are reported in Table 1. The inclusion of NNLO corrections leads to improvements, albeit small, across all sets of data with the exception of $pp\to hX$, where $\chi^{2}$ remains essentially unchanged.
@@
As expected, the total up and down quark PDFs are well determined. The gluon helicity PDF is seen to be clearly positive, with relatively small uncertainty, over a broad region in $x$ accessed by RHIC data. Furthermore, there is a pronounced pattern of positive $\Delta\bar{u}$ and negative $\Delta\bar{d}$ at moderately large $x$, while the strangeness helicity PDF is essentially undetermined. This is in contrast to our previous analyses [34, 2], and results from the fact that we no longer impose constraints related to SU(2) and SU(3) symmetry [^4] Nevertheless, the central NLO and NNLO fits both deviate only by a few per cent from these constraints that are usually given in terms of $F$ and $D$ values., as well as from the reduced constraints from SIDIS data due to the use of an $x$-cut. Sizable uncertainties remain for the sea quark and gluon PDFs at $x\lesssim 10^{-3}$, consistent with the lack of experimental data constraining that region.
Figures 35 show detailed comparisons of our fit results at NLO and NNLO accuracy to a few selected sets of $pp$ and SIDIS data that were used in our analysis. For comparison, we also show in Figs. 15 the results for an (unpublished) NLO fit (labeled as BDSSV22) obtained using the same methodology of [2], but updating the data sets to additionally include RHIC dijet and $W^{\pm}$ production data (for a collection of RHIC data, see, e.g. [43]), as well as COMPASS final SIDIS measurements [44]. As can be seen, the differences to our new NLO fit are rather moderate and mostly within the estimated uncertainty bands.
@@
Our study sets the stage for future NNLO global analyses of nucleon helicity structure in the EIC era. There are several avenues for future improvements, among them derivations of better approximations for the NNLO terms in $pp$ scattering, the implementation of the full NNLO results for SIDIS, detailed explorations of the residual scale dependence and power corrections and the inclusion of theoretical uncertainties in the fit [45, 46, 47].
Both the NLO and NNLO sets of replicas are publicly available in LHAPDF format in https://github.com/igborsa/BDSSV24.